# GranQ: Granular Zero-Shot Quantization with Unified Layer-Channel Awareness

### v1

@@ 3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table LABEL:tab6 in the supplementary material.
@@ 4.1 Observation
Furthermore, Figure 3 provides a visual representation of the effect of quantization on activation values across layers. In layer-quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 3 is provided in the supplementary material, in Figures LABEL:sup_fig3 and LABEL:sup_fig1.
@@ 4.2.2 Vectorized Quantization
This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors. A detailed computational process is provided in Algorithm LABEL:alg1, available in the supplementary material.
@@ 5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [27], and ImageNet (ILSVRC 2012) [11] validation datasets. For the CIFAR datasets, ResNet-20 [19] was used as the quantization model, whereas ResNet-18 [19], ResNet-50 [19], and MobileNetV2 [38] were employed for ImageNet. All experiments were conducted using the SGD optimizer [37] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [41, 10, 36, 5, 1, 14, 32, 23, 26]. For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35]. Further implementation details are provided in the supplementary material, Section LABEL:sec:Details_of_Experimental_Setup. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@ 5.3.2 Efficiency Evaluation
The first two methods are conventional quantization techniques: layer- and channel-wise. The third is scalar-based iterative, which considers all channels in each layer but performs multiple iterations. The fourth method is GranQ, which integrates both layer and channel information while employing vectorized quantization. Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3. In contrast, GranQ enhances efficiency through parallelized computations (Algorithm LABEL:alg1 in supplementary material).
@@ 6.1 Why GranQ is Effective?
The most common activation quantization approaches are known for their low computational cost and fast processing speeds. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 3). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section LABEL:sec:extraanalysis in the supplementary material). These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.

### v3

@@ 3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the supplementary material.
@@ 4.1 Observation
Furthermore, Figure 3 provides a visual representation of the effect of quantization on activation values across layers. In layer-quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 3 is provided in the supplementary material, in Figures 6 and 7.
@@ 4.2.2 Vectorized Quantization
This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors. A detailed computational process is provided in Algorithm 1, available in the supplementary material.
@@ 5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [27], and ImageNet (ILSVRC 2012) [11] validation datasets. For the CIFAR datasets, ResNet-20 [19] was used as the quantization model, whereas ResNet-18 [19], ResNet-50 [19], and MobileNetV2 [38] were employed for ImageNet. All experiments were conducted using the SGD optimizer [37] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [41, 10, 36, 5, 1, 14, 32, 23, 26]. For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35]. Further implementation details are provided in the supplementary material, Section B. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@ 5.3.2 Efficiency Evaluation
The first two methods are conventional quantization techniques: layer- and channel-wise. The third is scalar-based iterative, which considers all channels in each layer but performs multiple iterations. The fourth method is GranQ, which integrates both layer and channel information while employing vectorized quantization. Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3. In contrast, GranQ enhances efficiency through parallelized computations (Algorithm 1 in supplementary material).
@@ 6.1 Why GranQ is Effective?
The most common activation quantization approaches are known for their low computational cost and fast processing speeds. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 3). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section D in the supplementary material). These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.
@@ 7 Conclusion
Supplementary Material
Appendix A Notation
Symbol Definition
$x$ Original activation value
$x_{q}$ Quantized activation value
$x_{\text{deq}}$ Dequantized activation value
$x_{\min}$, $x_{\max}$ Minimum and maximum activation values
$Q$ Quantization operator
$s$ Scaling factor
$z$ Zero-point for quantization
$b$ Number of quantization bits
$\min\limits_{h,w}(x)$, $\max\limits_{h,w}(x)$ Channel-wise min/max activations
$\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
[…]
$\|\cdot\|_{2}$ L2 norm (Euclidean norm)
$\in$ Set membership (”element of”)
Mathematical Expressions
Forward Quantization $x_{q}=\left\lfloor\frac{x}{s}+z\right\rceil$
Dequantization $x_{\text{deq}}=s(x_{q}-z)$
Scaling Factor $s=\frac{x_{\max}-x_{\min}}{2^{b}-1}$
Zero-point $z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$
Cosine Similarity $\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}$
Relative Error $\frac{\|X-Q\|_{2}}{\|X\|_{2}}$
Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
In this section, we provide an overview of the main symbols and definitions employed throughout the paper. It specifically includes essential elements such as the original activation ($x$), quantized activation ($x_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
Appendix B Details of Experimental Setup
We provide a detailed description of the datasets, experimental setup, and comparison methods utilized in our experiments. This ensures the implementation of the experiments and clearly highlights the strengths of the proposed approach over existing methods.
B.1 Datasets
This paper outlines the experimental settings for CIFAR-10 [9], CIFAR-100 [9], and ImageNet (ILSVRC-2012) [4] datasets. CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images. ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments. We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
B.2 Experimental Setup
The data generation process followed the settings of AdaDFQ [14]. Specifically, ACGAN [12] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [14]. For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD [11] as the optimizer, with a momentum of 0.9 and a weight decay of 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch [13], and visualizations were generated with matplotlib [7]. All experiments were conducted on an NVIDIA RTX 3090 GPU.
B.3 Comparison Methods
GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ). Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
Algorithm 1 GranQ: Layer-Channel Quantization
1:Activation tensors $\{x_{l}\}_{l=1}^{L}$, where $x_{l}\in\mathbb{R}^{C\times(HW)}$, quantization bit $b$
2:Dequantized activation tensors $\{x_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do $\triangleright$ Process each layer separately
4:       Compute per-channel min/max values:
$\vec{x}_{\min}=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$
$\triangleright$ Compute along the channel axis for layer $l$
5:       Compute scaling factors and zero-points (Vectorized):
$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$
6:       Quantize values (Vectorized):
$x_{q,l}=\left\lfloor\frac{x_{l}}{\vec{s}}+\vec{z}\right\rceil$
$\triangleright$ Element-wise quantization for all channels in layer $l$
7:       Dequantize values (Vectorized):
$x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$
$\triangleright$ Reconstruct activation tensor for layer $l$
8:end for
9:Return the dequantized tensors:
$\textbf{return}\ \{x_{\text{deq},l}\}_{l=1}^{L}$
  • GDFQ [15]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation. It demonstrated that effective quantization is possible even without access to original data.
  • AIT [3]: AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation. It also provided a systematic analysis of training strategies in the ZSQ domain.
  • AdaDFQ [14]: AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
  • TexQ [2]: TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
  • RIS [1]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
  • PLF [5]: PLF was the first to evaluate generated data before quantization, offering a method to ensure more reliable data.
  • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
  • AKT [6]: AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
  • SynQ [8]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
Figure 6: Ablation study on relative error in ResNet-20 across layers. Layer-Q, Channel-Q, and GranQ represent layer-wise, channel-wise, and proposed quantization. Experiments used AdaDFQ on CIFAR-100 with 3-bit quantization.
Appendix C Algorithm
GranQ is a granular quantization method that processes activation values across all channels within each layer. Algorithm 1 describes the procedure for quantizing the activation tensor $\{x_{deq,l}\}^{L}_{l=1}$. The activation tensor of layer $l$ has a shape of $C\times(HW)$, and the minimum and maximum values are calculated independently for each channel. Based on these values, a scaling factor and zero-point are computed per channel for the quantization process. The resulting quantized values are dequantized back to floating-point representation prior to returning.
Appendix D Extra Analysis
D.1 Analysis of Activation Preservation
To analyze the severe quantization error (activation distortion) occurring in specific layers, we conducted additional experiments, as illustrated in Figure 6. This figure presents the relative error for each layer of ResNet-20 and reveals that certain layers such as the 18th exhibit significantly high errors. These findings suggest that both layer- and channel-wise quantization fail to effectively preserve activations in specific layers. In contrast, GranQ achieves consistently low errors across all layers, ensuring better activation preservation.
D.2 Analysis of Activation Magnitude
In Figure 7, we observe that channel-wise quantization is ineffective in ZSQ settings, similar to layer-wise quantization. As shown in Figure 7(a), the full-precision (FP) model exhibits significant activation magnitude differences across layers, even within the same channel axis. This highlights the limitations of channel-wise quantization in handling quantization errors. Conversely, GranQ addresses this issue by accounting for each channel across all layers, thereby minimizing quantization errors. This is particularly important in the final layers, where maintaining the activations of the FP model is critical for accurate inference. Whereas (a) fails to preserve these distributions, (b) successfully maintains them.
(a) Layer Quantization
(b) GranQ (Ours)
Figure 7: Visualization of activation magnitudes across layers in ResNet-20. (a) illustrates the activation magnitudes with layer-wise quantization, whereas (b) presents the results after applying the proposed GranQ method. In each subplot, the left side depicts activation magnitudes before quantization (32-bit), whereas the right side shows them after quantization (3-bit). The proposed GranQ effectively preserves the activation distribution, mitigating distortions caused by quantization.